OpenStudy (anonymous):

(3x^29y^23)^(-1/5) simplify

OpenStudy (anonymous):

ok do you take the 3 in the the problem and raise it to the (-1/5)?

OpenStudy (chaise):

You take the entire equation down and raise it to the POSITIVE 1/5 This can be written again as the 5th root of your equation. Let me do the entire thing in equation editor.

OpenStudy (chaise):

\[(3x^29y^23)^{-1/5}\] \[= 1/(3x^29y^23)^{1/5}\] \[= 1/\sqrt[5]{(3x^29y^{23}}^1\] Notice how the denominator in the equation goes into the square root and the numerator stays up?

OpenStudy (anonymous):

yes, is the simpliest form?